OpenStudy (anonymous):

y=log(base10)^3-4x? ....trying to figure out how to start?... any help is appreciated

OpenStudy (ranga):

Solve for x?

OpenStudy (anonymous):

I'm actually trying to just put it into exponential form so I can figure out the graph and find the domain. I guess I should of included that.

OpenStudy (ranga):

If \[\Large \log_{b} a = c\] then: \[\Large a = b ^{c}\]

OpenStudy (ranga):

\[\Large \log_{10} (3 - 4x) = y\]Therefore\[\Large (3 - 4x) = 10^{y}\]

OpenStudy (anonymous):

Ok. so 3-4x=10^x?

OpenStudy (anonymous):

Ohhhhh....:) ok. thank you. I super appreciate it.

OpenStudy (ranga):

^^^ the right side side should be 10^y

OpenStudy (ranga):

But to graph it you don't need to put it in an exponential form. You are already given y as a function of x and so you can find the domain and graph it without having to put it in exponential form.

OpenStudy (anonymous):

Oh. Well how does the 3-4x change the behavior of the graph does is reflect it across the x axis or is it still the same behavior as log b of x just as x>0 the line is increasing or decreasing?

OpenStudy (ranga):

First figure out the domain. The rule is you cannot take the logarithm of zero or a negative number. Can you figure out the domain here?

OpenStudy (anonymous):

oh. yes. I figured it out. I would set 3-4x>0 -3 -3 -4x>-3 /-4 /-4 x< 3/4 domain: (-infi.- 3/4)

OpenStudy (ranga):

Good. We know that x = 3/4 is a vertical asymptote for this function because when x approaches 3/4 we are approaching closer to log(0) which is negative infinity. So try a few values of x < 3/4 such as 0, -1, -2, etc. Knowing the general curve of a logarithmic function and knowing the vertical asymptote and a few points you can graph the function.

OpenStudy (anonymous):

great. thank you.

OpenStudy (ranga):

you are welcome.